OpenStudy (anonymous):

What is the volume of a square pyramid with base edges of 60 inches and a height of 10 inches? 36,000 in3 600 in3 12,000 in3 6,000 in3

jimthompson5910 (jim_thompson5910):

The area of the base is A = s^2 = 60^2 = 3600 square inches The volume is then Volume = (Area of base*height)/3 V = (Ah)/3 V = (3600*10)/3 V = 36000/3 V = 12000 So the volume is 12,000 cubic inches, which is choice C

OpenStudy (anonymous):

the numbers are different :c there & thanks !!